# Global Extraction of the $\rm{}^{12}C$ Nuclear Electromagnetic Response Functions (${\cal R}_{L}$ and ${\cal R}_{T}$) and Comparisons to Nuclear Theory and Neutrino/Electron Monte Carlo Generators

### v2

@@ II.5 Electron scattering at 180
where the $eff$ subscript denotes Coulomb corrected quantities described below.
II.6 ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region
As mentioned earlier, in the nuclear excitation region $4016<E_{x}<$ 40 MeV we use the precise (±$\pm 3\%$) Yamaguchi measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$ (where available). For $E_{x}<16$ MeV we extract ${\cal R}_{L}$ and ${\cal R}_{T}$ from fits Bodek and Christy (2023) to the nuclear excitation form factors, and performing a Gaussian smearing with the width determined by the experimental resolution. At $Q^{2}=0.01$ GeV2 we extract ${\cal R}_{T}$ from electron scattering cross sections at $E_{0}=55$ and 70 MeV and $\theta=180^{\circ}$ published in Goldemberg:64 Goldemberg and Barber (1964) and electron scattering cross sections at $E_{0}=65$ MeV and $\theta$ =180 published in Deforest:65 Deforest et al. (1965). At $Q^{2}=0$ we extract ${\cal R}_{T}$ from measured photo-absorption cross sections.
@@ III Coulomb Corrections
Including Coulomb corrections we define:
@@
$\epsilon^{CC}$ $=$ $\left[1+2\left(1+\frac{\nu^{2}}{Q^{2}_{eff}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
@@ IV Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$
From the Christy-Bodek overall fit we extract the relative normalization ($N_{i}$) of the various data sets as given in Table 2. For low excitation energy ($E_{x}<$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of excitation energy as described below. Similarly for ($E_{x}>$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of the square of the final state invariant mass $W^{2}$.
@@
We use the fit to apply $Q^{2}$ bin centering corrections ($C_{i}$) to all extracted value off $\Sigma_{i}^{\prime}$ that are within the $Q_{eff}^{2}$ range for each $Q_{eff}^{2}$ and of $W_{eff}^{2}$ bin such that they reflect the values at $Q_{enter}^{2}$, $W_{center}^{2}$, and $\epsilon_{center}$ for each cross section measurement,
$C_{i}$ $=$ $\frac{\epsilon{\cal R}^{fit}_{L}({Q_{center}^{2})}+\frac{1}{2}(\frac{{\bf q_{center}}}{Q_{center}})^{2}{\cal R}^{fit}_{T}({Q_{center}^{2}})}{\epsilon{\cal R}^{fit}_{L}({Q_{eff}^{2})}+\frac{1}{2}(\frac{{\bf q_{eff}}}{Q_{eff}})^{2}{\cal R}^{fit}_{T}({Q_{eff}^{2}})}.$
@@
After the application of the bin centering corrections to all extracted values $\Sigma^{\prime}_{i}$ in each $Q_{eff}^{2}$ bin we can assume that extracted values of $\Sigma_{i}$ are at $Q^{2}_{center}$. When we apply these corrections, we keep the final state invariant mass $W^{2}$ fixed for $\nu>$ 50 MeV, and only correct for the change in $Q_{eff}^{2}$. For $\nu<$ 50 MeV we use the excitation energy $E_{x}$ instead of $W^{2}$.
For $E_{x}$ bins at fixed values of $Q^{2}$ the value of $\nu$ in the bin center is extracted from value of $E_{x}$ in the bin center using the following expression.
$\nu$ $=$ $E_{x}+\frac{Q^{2}}{2M_{A}},$ (36)
where $M_{A}$ is the mass of the nuclear target (for carbon $M_{A}$=11.178 GeV). For $W^{2}$ bins at fixed values of $Q^{2}$, the value of $\nu$ at the bin center is extracted from value of $W^{2}$ in the bin center using the following expression:
$W^{2}$ $=$ $M^{2}+2M\nu-Q^{2},$ (37)
For $E_{x}$ bins at fixed values of $\bf q$ the value of $\nu$ at the bin center is extracted from value of $E_{x}$ in the bin center by solving the following quadratic equation:
$\nu$ $=$ $E_{x}+({\bf q}^{2}+\nu^{2})/(2M_{A})$ (38)
For $W^{2}$ bins at fixed values of $\bf q$ the value of $\nu$ in the bin center is extracted from value of $W^{2}$ in the bin center by solving the following quadratic equation:
$W^{2}$ $=$ $M^{2}+2M\nu-({\bf q}^{2}+\nu^{2}).$ (39)
IV.1 Comparison to Previous Measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
@@
A comparison between our extraction of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ (red diamonds) to previous extractions are shown in Fig. 7 for the three ${\bf q}$ values used in Jourdan:96 Jourdan (1996, 1995). In the Barreau:82 Barreau et al. (1983a) analysis the response functions were extracted for only three $\bf q$ values (0.30, 0.40 and 0.55 GeV) using only their own (Saclay) cross section measurements. A later extraction by Jourdan:96 was also performed for three slightly different values of momentum transfer (${\bf q}=0.30$, 0.38 and 0.57 GeV). In the Jourdan analysis the same Barreau:82 cross sections used but with additional cross section data from SLAC.
@@
There is good agreement between our extraction at $Q^{2}$=0.093 GeV2222 to the extraction by Baran:1988 Baran et al. (1988) at $Q^{2}$=0.1 GeV2 (light green circles) in the region of the $\Delta$(1238) resonance as seen in Figure 7(B).
@@ V Comparison to theoretical predictions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ extracted from the Christy-Bodek universal fit (which includes nuclear excitations) are shown as solid black lines, and the QE contributions to the fit (including TE) are the dashed black lines.
@@
    “Green’s Function Monte Carlo” (GFMC)Lovato et al. (2016) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states. GFMC predictions are only available for $0.3\leq{\bf q}\leq 0.57$ GeV.
@@
    “Short Time Approximation Quantum Monte Carlo” (STA-QMC)Andreoli et al. (2024) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states (for both electron and neutrino scattering processes). Currently, the STA-QMC predictions are only valid for $0.3\leq{\bf q}\leq 0.76$ GeV.
@@
The improved SuSAv2 predictions for the 1p1h and 2p2h channels has been implemented in genie  for electrons and neutrinos Dolan et al. (2020); Papadopoulou et al. (2021), and implementation in neut is underway. In addition, the SuSAv2-inelastic model is currently being implemented in genie for electrons and neutrinos. The SuSaA2 2p2h channel has been implemented (only for neutrinos) in NuWro,  andSTA-QMC been implementedBarrow et al. (2021) in genie  but only for $\rm He^{4}$ (though an hadron tensor interfaceGardiner (2019)), and the achilles generator is being developed for use in future neutrino experiments.
@@ V.1 Comparison to the predictions of GFMC, STA-QMC, and ED-RMF
The extracted ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are compared to predictions of GFMC in Fig. 9, STA-QMC in Fig. 9-10) and ED-RMF in Fig. 8-10). The three theoretical calculations include contributions from both 1-body and 2-body (1b+2b) currents. In these three calculations there is enhancement of the calculated transverse response function ${\cal R}_{T}(\bf q,\nu)$ The transverse enhancement (TE) primarily originates from the interference Simons et al. (2022) between 1b and 2b currents which result in 1p1h. (1 nucleon 1 hole) final states. The contribution of 2b currents to ${\cal R}_{T}(\bf q,\nu)$ which results in 2p2h (2 nucleons 2 holes) final states are not included ED-RMF but are included in GFMC and STA-QMC. Note that final states with pions are not included in any of the three calculations (pion production is only included in the improved version of SuSAv20).
The GFMC predictions are shown as the pink lines in Fig. 9. These calculations are rather complex and require a large number of CPU hours, especially at low $\bf q$. Consequently the response functions below $\bf q$=0.3 GeV and above $\bf q$=0.57 are not available for GFMC.
The predictions for $\rm{}^{12}C$ (shown as the light green lines in Fig. 9-10) are provided for $0.3\leq{\bf q}\leq 0.76$ GeV. By its nature STA-QMC only works for $\bf q>$ 0.3 GeV. The STA-QMC predictions for for $0.3\leq{\bf q}\leq 0.6$ GeV are in agreement with our measurements. Presently, STA-QMC does not fully account for relativistic effects at the vertex. Therefore, although STA-QMC calculations have been provided for $\bf q$=0.76 GeV (shown in Fig. 10), it is expected that that they will be less accurate in the high $\bf q$ region where relativistic effects become increasingly more relevant. Although the STA-QMC calculations are based on the same many-nucleon Hamiltonian and electromagnetic currents as GFMC, the STA-QMC calculations are in better agreement with our measurements than GFMC.
The ED-RMF theoretical predictions are shown as the blue lines. These predictions (which are available at all values of $\bf q$) include contributions to 1p1h final states from QE scattering as well as from nuclear excitations. At higher values of $\bf q$ and higher values of $\nu$ the ED-RMF predictions for ${\cal R}_{T}(\bf q,\nu)$ are below the data. This is expected because the ED-RMF theoretical calculations for ${\cal R}_{T}(\bf q,\nu)$ do not include 2-body current processes with 2p2h final states or Fermi smeared pion production processes. The STA-QMC calculation includes both single and two nucleon final states but is only valid over a more restricted kinematic range and does not account for nuclear excitations. The calculations of both ED-RMF and STA-QMC are directly applicable to the same kinematic regions for neutrino scattering.
@@ V.1.1 The ED-RMF theoretical formalism
For the final state, the momentum distributions for the particle states in the continuum are computed by solving the Dirac equation with the energy-dependent relativistic mean-field (ED-RMF) potential, which has no imaginary part so no flux is lost. The ED-RMF is the same potential used in the bound state but multiplied by a function that weakens it for increasing energy (for see details ref. González-Jiménez et al. (2019); Nikolakopoulos et al. (2019)). At low energies of the final nucleon (up to about 20 MeV in the continuum) the ED-RMF potential is the same potential employed to compute the momentum distributions for the initial bound nucleons an at larger energies it approaches the behavior of a phenomenological optical potentials. This ensures orthogonality between the initial and final states at these low energies and thus the reliability of the cross-section computed in this region. In this way, the ED-RMF calculation also preserves the contribution (and resonant character) from single-channel resonances corresponding to excitations appearing slightly above the single-nucleon emission threshold of the 12C final system, which subsequently decay by single nucleon emission.

### v3

@@ II.4 Electron scattering at 180
where the effeff subscript denotes Coulomb corrected quantities described below.
II.5 ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region
As mentioned earlier, in the nuclear excitation region $16<E_{x}<40$ MeV we use the precise ($\pm 3\%$) Yamaguchi measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$ (where available). For $E_{x}<16$ MeV we extract ${\cal R}_{L}$ and ${\cal R}_{T}$ from fits Bodek and Christy (2023) to the nuclear excitation form factors, and performing a Gaussian smearing with the width determined by the experimental resolution. At $Q^{2}=0.01$ GeV2 we extract ${\cal R}_{T}$ from electron scattering cross sections at $E_{0}=55$ and 70 MeV and $\theta=180^{\circ}$ published in Goldemberg:64 Goldemberg and Barber (1964) and electron scattering cross sections at $E_{0}=65$ MeV and $\theta$ =180 published in Deforest:65 Deforest et al. (1965). At $Q^{2}=0$ we extract ${\cal R}_{T}$ from measured photo-absorption cross sections.
@@ III Coulomb Corrections
Including Coulomb corrections (CC) we define:
@@
$E_{x}^{cc}$ $=$ $E_{x}$
$W^{2}_{A-eff}$ $=$ $M^{2}_{A}+2M_{A}\nu-Q^{2}_{eff}$
$\epsilon^{CC}$ $=$ $\left[1+2\left(1+\frac{\nu^{2}}{Q_{eff}^{2}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
@@ IV Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$
From the Christy-Bodek 2024 overall fit we extract the relative normalization ($N_{i}$) of the various data sets as given in Table 2. For low excitation energy ($E_{x}<$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of the final state excited nuclear mass $W^{2}_{A}$, as described below. Similarly for ($E_{x}>$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of the square of the final state mass defined in terms of the mass of the nucleon $W^{2}$.
@@
We use the fit to apply $Q^{2}$ bin centering corrections ($C_{i}$) to all extracted value of $\Sigma_{i}^{\prime}$ that are within the $Q_{eff}^{2}$ range for each $Q_{eff}^{2}$ and of $W_{eff}^{2}$ bin such that they reflect the values at $Q_{enter}^{2}$, $W_{center}^{2}$, and $\epsilon_{center}$ for each cross section measurement,
IV.1 Bin centering corrections for fix $Q^{2}$ bins
$C_{i}=\frac{\epsilon_{center}{\cal R}^{fit}_{L}+\frac{1}{2}(\frac{{\bf q_{center}}}{Q_{center}})^{2}{\cal R}^{fit}_{T}}{\epsilon^{cc}{\cal R}^{fit}_{L}(W^{2}_{eff},{Q_{eff}^{2})}+\frac{1}{2}(\frac{{\bf q_{eff}}}{Q_{eff}})^{2}{\cal R}^{fit}_{T}({W^{2}_{eff},Q_{eff}^{2}})},$
For excitation energies above 50 MeV we use $W^{2}$ based bin centering corrections.
${\cal R}^{fit}_{L}={\cal R}^{fit}_{L}(W^{2}_{center},{Q_{center}^{2})}$
${\cal R}^{fit}_{T}={\cal R}^{fit}_{T}(W^{2}_{center},{Q_{center}^{2})}$
$\nu_{center}=\frac{W^{2}_{center}-M^{2}+Q^{2}_{center}}{2M}$ (37)
and
$\epsilon_{center}=\left[1+2\left(1+\frac{\nu^{2}_{center}}{Q_{center}^{2}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
@@ IV.1 Bin centering corrections for fix $Q^{2}$ bins
After the application of the bin centering corrections to all extracted values $\Sigma^{\prime}_{i}$ in each $Q_{eff}^{2}$ bin we can assume that extracted values of $\Sigma_{i}$ are at $Q^{2}_{center}$. When we apply these corrections, we keep the final state mass $W^{2}$ fixed for $\nu>$ 50 MeV, and only correct for the change resulting from changing $Q_{eff}^{2}$ to $Q^{2}_{center}$, which also results in changing $\nu$ to $\nu_{center}$ and $\epsilon_{cc}$ to $\epsilon_{center}$.
For $Ex<$ 50 MeV at fixed values of $Q^{2}$ the value of $\nu_{center}$ is
$\nu_{center}=E_{x}+\frac{Q^{2}_{center}}{2M_{A}}$ (39)
where $M_{A}$ is the mass of the nuclear target (11.178 GeV for carbon). and
$\epsilon_{center}=\left[1+2\left(1+\frac{\nu^{2}_{center}}{Q_{center}^{2}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
$W^{2}_{center}=M^{2}+2M\nu_{center}-Q^{2}_{center}$
${\cal R}^{fit}_{L}={\cal R}^{fit}_{L}(W^{2}_{center},{Q_{center}^{2})}$
${\cal R}^{fit}_{T}={\cal R}^{fit}_{T}(W^{2}_{center},{Q_{center}^{2})}.$ (40)
For fixed $Q^{2}$ bins, after bin centering, all of the Rosenbluth fits are done as a function of of $\epsilon_{center}$, and the extracted ${\cal R}_{L}$ and ${\cal R}_{T}$ are plotted as a function of $\nu_{center}$.
IV.2 Bin centering corrections for fixed $\bf q$ bins
For $E_{x}>$ 50 MeV. we solve the following quadratic equation for $\nu_{center}$.
$W^{2}_{center}$ $=$ $M^{2}+2M\nu_{center}-({\bf q}^{2}_{center}-\nu^{2}_{center}),$
which yields
$Q^{2}_{center}={\bf q}^{2}_{center}-\nu^{2}_{center,}$
and then follow up with equations IV.1 and IV.1.
For $Ex<$ 50 MeV at fixed $\bf q$ beins We solve the following quadratic equation for $\nu_{center}$.
$\nu_{center}$ $=$ $E_{x}+\frac{{\bf q}^{2}-\nu^{2}_{center}}{2M_{A}}$ (41)
which yields
$Q^{2}_{center}={\bf q}^{2}_{center}-\nu^{2}_{center},$
$W^{2}_{center}=M^{2}+2M\nu_{center}-Q^{2}_{center},$
and then follow up with equation IV.1.
IV.3 Comparison to Previous Measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
@@
A comparison between our extraction of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ (red diamonds) to previous extractions are shown in Fig. 7 for the three $\bf q$} values used in Jourdan:96 Jourdan (1996, 1995). In the Barreau:82 Barreau et al. (1983a) analysis the response functions were extracted for only three $\bf q$ values (0.30, 0.40 and 0.55 GeV) using only their own (Saclay) cross section measurements. A later extraction by Jourdan:96 was also performed for three slightly different values of momentum transfer (${\bf q}=0.30$, 0.38 and 0.57 GeV). In the Jourdan analysis the same Barreau:82 cross sections used, with additional cross section data from SLAC.
@@
There is good agreement between our extraction at $Q^{2}$=0.093 GeV22 and the extraction by Baran:1988 Baran et al. (1988) at $Q^{2}$=0.1 GeV2 (light green circles) in the region of the $\Delta$(1238) resonance as seen in Figure 7(B).
IV.4 Systematic Errors
We estimate an overall systematic error of 4% in the extracted values of the longitudinal and transverse response functions. This error is anti-correlated between ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ .
@@ V Comparison to theoretical predictions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ extracted from the Christy-Bodek 2024 universal fit (which includes nuclear excitations) are shown as solid black lines, and the QE contributions to the fit (including TE) are the dashed black lines.
@@
    “Green’s Function Monte Carlo” (GFMC)Lovato et al. (2016, 2023a) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states. GFMC predictions are only available for $0.3\leq{\bf q}\leq 0.57$ GeV.
@@
    “Short Time Approximation Quantum Monte Carlo” (STA-QMC)Andreoli et al. (2024) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states (for both electron and neutrino scattering processes). Currently, the STA-QMC predictions are valid for $0.3\leq{\bf q}\leq 0.65$ GeV.
@@
The improved SuSAv2 predictions for the 1p1h and 2p2h channels has been implemented in genie  for electrons and neutrinos Dolan et al. (2020); Papadopoulou et al. (2021), and implementation in neut is underway. In addition, the SuSAv2-inelastic model is currently being implemented in genie for electrons and neutrinos. The SuSAv2 2p2h channel has been implemented (only for neutrinos) in NuWro,  the achilles generator is being developed for use in future neutrino experiments, and STA-QMC been implementedBarrow et al. (2021) in genie  for $\rm He^{4}$ (though an hadron tensor interfaceGardiner (2019)).
@@ V.1 Comparison to the predictions of GFMC, STA-QMC, and ED-RMF
The extracted ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are compared to predictions of GFMC in Fig. 9, STA-QMC in Fig. 9-10) and ED-RMF in Fig. 8-10). The three theoretical calculations include contributions from both 1-body and 2-body (1b+2b) currents. In these three calculations there is enhancement of the calculated transverse response function ${\cal R}_{T}(\bf q,\nu)$ . The transverse enhancement (TE) primarily originates from the interference Carlson et al. (2002); Simons et al. (2022); Andreoli et al. (2024) between 1b and 2b currents which result in 1p1h. (1 nucleon 1 hole) final states. The contribution of 2b currents to ${\cal R}_{T}(\bf q,\nu)$ which results in 2p2h (2 nucleons 2 holes) final states is not included ED-RMF but are included in GFMC and STA-QMC. Note that final states with pions is not included in any of the three calculations (pion production is only included in the improved version of SuSAv2).
The GFMC predictions are shown as the pink lines in Fig. 9. These calculations are rather complex and require a large number of CPU hours, especially at low $\bf q$. Consequently the response functions below $\bf q$=0.3 GeV and above $\bf q$=0.57 are not available for GFMC. In addition, there are large uncertainties in the predictions near threshold (small $\nu$).
The STA-QMC predictions for $\rm{}^{12}C$ (shown as the light green lines in Fig. 9-10) are provided for $0.3\leq{\bf q}\leq 0.76$ GeV. By its nature STA-QMC only works for $\bf q>$ 0.3 GeV. The STA-QMC predictions for $0.3\leq{\bf q}\leq 0.6$ GeV are in agreement with our measurements. Presently, STA-QMC does not fully account for relativistic effects at the vertex. Therefore, although STA-QMC calculations have been provided for up to $\bf q$=0.76 GeV (shown in Fig. 10), it is expected that that they will be less accurate in the high $\bf q$ region where relativistic effects become increasingly more relevant. The STA-QMC calculations are based on the same many-nucleon Hamiltonian and electromagnetic currents as GFMC. Since STA-QMC does not have any knowledge of the correct threshold behavior or low-energy properties of the system, the correct threshold behavior is imposed as a constraintPastore et al. (2020).
The ED-RMF theoretical predictions are shown as the blue lines. These predictions (which are available at all values of $\bf q$) include contributions to 1p1h final states from QE scattering as well as from nuclear excitations. At higher values of $\bf q$ and higher values of $\nu$ the ED-RMF predictions for ${\cal R}_{T}(\bf q,\nu)$ are below the data. This is expected because the ED-RMF theoretical calculations for ${\cal R}_{T}(\bf q,\nu)$ do not include 2-body current processes with 2p2h final states or Fermi smeared pion production processes. The STA-QMC calculation includes both single and two nucleon final states but is valid over a more restricted kinematic range and does not account for nuclear excitations. The calculations of both ED-RMF and STA-QMC are directly applicable to the same kinematic regions for neutrino scattering.
@@ V.1.1 The ED-RMF theoretical formalism
For the final state, the momentum distributions for the particle states in the continuum are computed by solving the Dirac equation with the energy-dependent relativistic mean-field (ED-RMF) potential, which has no imaginary part so no flux is lost. The ED-RMF is the same potential used in the bound state but multiplied by a function that weakens it for increasing energy (for see details ref. González-Jiménez et al. (2019); Nikolakopoulos et al. (2019)). At low energies of the final nucleon (up to about 20 MeV in the continuum) the ED-RMF potential is the same potential employed to compute the momentum distributions for the initial bound nucleons and at larger energies it approaches the behavior of a phenomenological optical potentials. This ensures orthogonality between the initial and final states at these low energies and thus the reliability of the cross-section computed in this region. In this way, the ED-RMF calculation also preserves the contribution (and resonant character) from single-channel resonances corresponding to excitations appearing slightly above the single-nucleon emission threshold of the 12C final system, which subsequently decay by single nucleon emission.